The measure of the angle is 81 degrees more than twice that of its supplement.
I have:
x+(2x+81)=180
3x=99
x=33
And now I'm stuck.

Answer by Alan3354(69443)

The measure of the angle is 81 degrees more than twice that of its supplement.
I have:
x+(2x+81)=180
----------
x = the angle
180 - x = its supplement
x = 2*(180 - x) + 81
x = 360 - 2x + 81 = -2x + 441
3x = 441
x = 147